# arXiv:2603.01334v1[cs.NI] 02 Mar 2026

### v1

@@ IV-A Static Techniques
A simple way to improve the energy efficiency of basic CGR is by applying thresholds based on the cloud cover forecast. Given a cloud cover threshold value of $T$ and a current cloud cover percentage $CC$, the scheme logic can be described as follows:
@@ IV-B Adaptive Techniques
where $\mathsf{CC}$ is the cloud cover for the next contact, $DV$ is the remaining data volume to be delivered, $\mathsf{CVR}$ is the remaining system capacity. $\mathsf{CVR}$ can be defined as follows:
@@
Finally, $SGC$ is a matrix containing all future weather and contact information for the downlink period. This matrix is an N by two matrix shown in the equation below:
@@
The delivery ratio and the total utilized contact time for an episode are defined as $DR$ and $CT$, respectively. With the scaling factor $c$, the terminal reward is defined as
@@
Given the discrete nature of the action space, value-based RL and DRL algorithms were selected. The two algorithms selected were Q-learning and DDQN. For the aforementioned algorithms, the observation information provided to both algorithms must include the future contact information $SGC$, which must be updated at each step. This involves removing from the $SGC$ the contacts that are no longer available and reorganizing the matrix to reflect the future contacts. This is due to the online nature of the problem. In addition, for the Q-learning algorithm, due to the continuous nature of the observation space, discretization must be conducted. This reduces the granularity of the information provided to the agent, but it is necessary due to the tabular nature of the algorithm. Q-learning was selected primarily as a point of comparison for the DDQN implementation. This was to ensure the added complexity of DDQN is justified by increased performance when compared to simple tabular RL approaches.
@@ II Number of Inputs
The time complexity of each algorithm is defined in terms of the number of contacts $N$ in each downlink period. For CGR and threshold schemes, the relationship between time complexity and $N$ is linear under worst-case assumptions. In addition, the constant $c_{LA}$ incorporates the time complexity of the link availability calculations, resulting in the following expressions:
@@
In terms of Algorithm 4, in addition to the linear relationship between $N$ and the number of times all the processes must be run, the sorting algorithm must sort the available contacts using a standardized sorting algorithm (assumed to be merge sort for this analysis). The complexity of sorting is defined as
@@
Unlike Algorithm 4, Algorithm 5 must conduct sorting operations every time a contact is used. In the worst-case scenario, this would entail conducting sorting operations for every contact $N$. Hence, the overall time complexity can be defined as follows:
@@
where $c_{\mathsf{FP}}$ is the number of calculations required for a single input and $I$ is the number of inputs. The number of contacts in the downlink period only affects the number of operations in an input neural network layer. The inference time complexity is linearly proportional to the number of contacts in a downlink period, and hence can be defined as
@@ V-A General Simulation Configurations
For the two general simulations, several key simulation parameters were set to be constant. The number of contacts $N$ in each downlink period was set to 10, and the volume of each contact $i$ was set to be of equal length. This more simplistic simulation configuration allows for an in-depth analysis of each scheme in a highly controlled environment.

### v2

# Energy Efficient Traffic Scheduling For Optical LEO Satellite Downlinks

@@ IV-A Static Techniques
A simple way to improve the energy efficiency of basic CGR is by applying thresholds based on the cloud cover forecast. Given a cloud cover threshold value of $T$ and a current cloud cover percentage CCCC, the scheme logic can be described as follows:
@@ IV-B Adaptive Techniques
where $\mathsf{CC}$ is the cloud cover for the next contact, DVDV is the remaining data volume to be delivered, $\mathsf{CVR}$ is the remaining system capacity. $\mathsf{CVR}$ can be defined as follows:
@@
Finally, SGCSGC is a matrix containing all future weather and contact information for the downlink period. This matrix is an N by two matrix shown in the equation below:
@@
The delivery ratio and the total utilized contact time for an episode are defined as DRDR and CTCT, respectively. With the scaling factor $c$, the terminal reward is defined as
@@
Given the discrete nature of the action space, value-based RL and DRL algorithms were selected. The two algorithms selected were Q-learning and DDQN. For the aforementioned algorithms, the observation information provided to both algorithms must include the future contact information SGCSGC, which must be updated at each step. This involves removing from the SGCSGC the contacts that are no longer available and reorganizing the matrix to reflect the future contacts. This is due to the online nature of the problem. In addition, for the Q-learning algorithm, due to the continuous nature of the observation space, discretization must be conducted. This reduces the granularity of the information provided to the agent, but it is necessary due to the tabular nature of the algorithm. Q-learning was selected primarily as a point of comparison for the DDQN implementation. This was to ensure the added complexity of DDQN is justified by increased performance when compared to simple tabular RL approaches.
@@ II Number of Inputs
The time complexity of each algorithm is defined in terms of the number of contacts NN in each downlink period. For CGR and threshold schemes, the relationship between time complexity and NN is linear under worst-case assumptions. In addition, the constant $c_{LA}$ incorporates the time complexity of the link availability calculations, resulting in the following expressions:
@@
In terms of Algorithm 4, in addition to the linear relationship between NN and the number of times all the processes must be run, the sorting algorithm must sort the available contacts using a standardized sorting algorithm (assumed to be merge sort for this analysis). The complexity of sorting is defined as
@@
Unlike Algorithm 4, Algorithm 5 must conduct sorting operations every time a contact is used. In the worst-case scenario, this would entail conducting sorting operations for every contact NN. Hence, the overall time complexity can be defined as follows:
@@
where $c_{\mathsf{FP}}$ is the number of calculations required for a single input and II is the number of inputs. The number of contacts in the downlink period only affects the number of operations in an input neural network layer. The inference time complexity is linearly proportional to the number of contacts in a downlink period, and hence can be defined as
@@ V-A General Simulation Configurations
For the two general simulations, several key simulation parameters were set to be constant. The number of contacts NN in each downlink period was set to 10, and the volume of each contact $i$ was set to be of equal length. This more simplistic simulation configuration allows for an in-depth analysis of each scheme in a highly controlled environment.